# DoubleAgents: Interactive Simulations for Alignment in Agentic AI

### v2

@@ Abstract.
Agentic workflows promise efficiency, but adoption hinges on whether people can align systems that act on their behalf with their goals, values, and situational expectations. We present DoubleAgents, an agentic planning tool that embeds transparency and control through user intervention, value-reflecting policies, rich state visualizations, and uncertainty flagging for human coordination tasks. A built-in respondent simulation generates realistic scenarios, allowing users to rehearse, refine policies, and calibrate their use of agentic behavior before live deployment. We evaluate DoubleAgents in a two-day lab study (n=10), three deployment studies, and a technical evaluation. Results show that participants initially hesitated to delegate but used simulation to probe system behavior and adjust policies, gradually increasing delegation as agent actions became better aligned with their intentions and context. Deployment results demonstrate DoubleAgents’ real-world relevance and usefulness, showing that simulation helps users effectively manage real-world tasks with higher complexity and uncertainty. We contribute interactive simulation as a practical pathway for users to iteratively align and calibrate agentic systems.
generative AI, AI agents, trust, reliance, transparency, state tracking, seminar organizing, control, planning, human coordination, reasoning, constitutional AI
copyright: acmcopyrightjournalyear: 2025conference: 2026; 2026; arxivbooktitle: Proceedings of the 2026price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
Figure 1. DoubleAgents is an agentic AI coordination tool that uses interactive simulation to let users experience and align the full ReAct cycle before live deployment. Starting from an initial setup (left)—including a simulation goal, simulated test-case personas, and an initial set of policies, and email templates—users can undergo multiple simulated ReAct cycles (middle) to rehearse planning, review and approve actions, execute plans, detect new edge cases, and iteratively refine policies via stop hooks, and email templates over repeated test cases and time. This process results in (right) the final speaker-slot allocation, as well as a user-aligned set of policies, emails, and stop hooks before the system sends real emails.
@@ 1. Introduction
Human–AI alignment is one of the defining challenges of the 21st century, especially for agentic AI: systems that go beyond answering questions to taking actions on users’ behalf (Leike et al., 2018; Soares and Fallenstein, 2014; Hendrycks et al., 2021). Following the Bidirectional Human–AI Alignment framework defined by Shen et al. (2025), we view alignment in agentic settings as an ongoing co-adaptation process between users and systems. On the system side, agents must be continuously shaped through interaction history and curated context so their actions reflect an individual user’s goals, preferences, and social norms (Zhang et al., 2025; Gao et al., 2024). On the human side, users must continually develop appropriate mental models, usage patterns, and AI literacy so they can effectively direct, evaluate, and calibrate delegation (Vanderlyn et al., 2025; Andrews et al., 2023). As agentic AI becomes more common, we need practical ways to assess and improve whether these systems remain aligned as they act in the world.
Today, many popular agentic use cases, such as agentic coding or “vibe coding”, are relatively low-stakes and operate in closed sandbox environments (Horikawa et al., 2025). In these domains, alignment is often easier because systems can leverage large-scale, well-structured public data (e.g., GitHub), and success criteria are comparatively legible (e.g., whether code compiles, passes tests, or produces the desired output) (Chen et al., 2021; Jimenez et al., 2024). However, in more complex tasks that must adhere to human and domain knowledge as well as social cognition and norms, such as writing emails to real people to request favors, where etiquette is highly context-dependent, agents cannot reliably align on their own (Afroogh et al., 2024; Gabriel, 2020). These tasks often require nuanced, case-by-case judgment that is difficult to infer from generic training data alone and may demand substantial human oversight if deployed naively.
One representative scenario is human coordination, such as scheduling workers with the right skills and experience for shifts they are available for or inviting researchers to a speaker series. In this work, we focus on the speaker-series scheduling setting. Organizers must secure a speaker for each slot (e.g., 4-5pm on Fridays), coordinate availability across multiple candidates, and handle uneven responsiveness (e.g., some invitees reply late or not at all). Responses can also introduce socially sensitive and context-dependent requests: “Can you reimburse travel?”, “Can I do Zoom?”, “I can’t make it on Fridays, what about Thursdays 4-5pm?” “Sorry for the last-minute cancellation—I now have a conflict,” or “My schedule is uncertain; can you follow up next month?” Determining the right next step is rarely optimal in a mathematical sense; it depends on the organizer’s values and preferences, which we refer to as “policies” (e.g., “Offer Zoom only if in-person is not feasible and no one else can do that slot,” “Prioritize senior faculty over PhD students when conflicts happen,” or “Follow up after two business days unless urgent”). In practice, because these “actions” are carried out almost entirely through email communication, the organizer must decide which email to send next (and to whom), how to phrase each email with the right content and right tone, and when to send it.
While an LLM-based agent may generate plausible actions and write acceptable emails, it often lacks the organizer’s specific norms and cannot reliably anticipate how preferences should change across cases. Alignment therefore must happen on both sides: the system needs ways to acquire and apply three user-specific automation refinements: (1) coordination policies that guide planning and action choices, (2) email templates and writing style that match the user’s voice and social intent for each action, and (3) stop hooks that define when the agents need to pause and request for user confirmation; and the user needs support to form a clear mental model and specify these three preferences, especially early on, when they may not yet know what rules, tone, or escalation boundaries they want. The central challenge lies in how can we accelerate this bidirectional alignment so the agent can act more autonomously over time, without requiring constant babysitting.
Our core approach is interactive simulation to support iterative automation refinement, enabling alignment between users and agents. As modern AI can simulate diverse human behaviors (Park et al., 2023, 2022), we use simulation as a testbed to expose users to realistic coordination scenarios, including problematic or edge-case behaviors, and to help them define and refine their communication strategies. During simulations, the agent proposes plans and actions (e.g., inviting multiple candidates, deciding when to wait versus follow up, or holding a slot for a high-priority speaker), and the user evaluates whether these choices are socially and contextually appropriate. This mirrors ideas from software engineering such as unit and integration testing: a system earns confidence not by succeeding on a single happy path, but by handling a range of representative cases. In our context, however, completing a task is not enough: agentic AI systems must be user-specific and value-aligned, reflecting personal priorities, norms, and preferences.
We present DoubleAgents, a system for aligning agentic AI through interactive simulation. DoubleAgents integrates generative AI to simulate two complementary roles, as Figure 2 shows. First, it uses a ReAct loop (Yao et al., 2023) that pairs coordination agents for planning and execution. The agents can call tools to retrieve relevant context, query the current task state, and execute actions such as drafting and sending emails to potential speakers. We extend the loop with explicit interaction checkpoints so users can review, intervene, and approve decisions throughout planning and execution. Second, DoubleAgents includes an LLM-based simulator of respondents that generates diverse, realistic replies for different personas (e.g., a professor who responds slowly or a respondent who introduces funding or logistics constraints). This simulation enables repeated run-throughs of difficult cases, helping users understand how the agent behaves, stress-test policies against edge cases, and iteratively refine preferences and autonomy boundaries. Across interactive simulation runs, the primary outputs are not only a candidate schedule (speaker–slot allocation), but also a reusable set of user-specific alignment artifacts: refined policies, updated email templates, and calibrated stop hooks that can be applied when the system is deployed to send real emails.
We evaluate DoubleAgents through a technical evaluation, a two-day lab study (n=10) in which participants used DoubleAgents with simulated respondents, and three real-world deployments. In the lab study, participants initially hesitated to delegate. However, over time, through multiple simulation rounds and the two study sessions, participants increasingly delegated, as visible state, policy-grounded reasoning, and clear intervention points (including stop-hook flagging) helped them calibrate autonomy and build confidence in the agent’s planning and execution. By the end, participants also produced reusable policies, stop hooks, and email templates aligned with their voice and style. Three week-long deployment studies demonstrate DoubleAgents’ usefulness and relevance for real-world, complex human coordination tasks. DoubleAgents was connected to organizers’ real email workflows and sending emails to real invitees, as well as incorporated incoming real speakers’ replies into its ongoing coordination state. Organizers reported that interactive simulation beforehand helped them prepare for deployment by crafting strong initial policies and email templates for real-world use.
Our work demonstrates how agentic systems can support reliable, human-centered coordination by embedding simulation, transparency, and interactive oversight into the agentic AI workflow. Specifically, we contribute:
  • A design perspective on alignment in agentic AI, realized through a policy-driven, intervention-friendly ReAct framework that supports transparency, human oversight, and alignment with human goals and values.
  • An interactive simulation mechanism that models realistic human responses, enabling users to safely explore agent behavior, understand system dynamics, rehearse interventions, and iteratively refine policies, email references, and stop hooks before real-world application.
  • Empirical results from a user study (n=10) and three real-world deployments, examining how interactive simulation and policy refinement support alignment in agentic AI for human coordination.
@@ 2.1. Alignment in Human-AI Systems
As AI agents take on more complex tasks, ensuring they align with human values has become a central research challenge. A key strategy is reinforcement learning from human feedback (RLHF), which fine-tunes models to produce outputs consistent with user expectations (Ouyang et al., 2022). However, alignment doesn’t end at training. Studies show that users often push back on AI responses that conflict with their values, steering behavior through real-time feedback, a process termed user-driven value alignment (Fan et al., 2025). Such interactive alignment is critical in high-stakes contexts like medical decision support, where LLM-powered co-pilot systems have helped clinicians catch more errors than either humans or AI alone (Ong et al., 2025). These findings show the importance of transparency, oversight, and human agency in AI deployment.
@@
Another approach, Constitutional AI, replaces constant human feedback with rule-based self-supervision. Models are trained to revise outputs using explicit principles like “Be respectful” (Bai et al., 2022). Recent work has expanded this into participatory frameworks such as Collective Constitutional AI, where large groups co-author alignment rules to reflect diverse values (Huang et al., 2024). Tools like ConstitutionMaker further help users turn natural language feedback into usable policies (Petridis et al., 2024). DoubleAgents contributes to this space by embedding alignment into interactive simulation. Users iteratively refine coordination policies, email tone, and stop hooks through rehearsed ReAct loops, treating alignment not as a one-time step, but as a continuous, participatory process.
2.2. Agentic AI in Task Planning and Execution
AI agents, systems that perform tasks on users’ behalf, operate through tool use (e.g., APIs) (Schick et al., 2023) or direct interface manipulation (Anthropic, 2025; OpenAI, 2025; Manus AI, 2025; Labs, 2025). While these agents show growing capabilities, delegating control to them still raises concerns around trust, transparency, and alignment (Afroogh et al., 2024). Planning and execution remain core challenges. Benchmarks like Tau-Bench, NaturalPlan, TravelPlanner, and AppWorld (Yao et al., 2024; Zheng et al., 2024; Xie et al., 2024; Trivedi et al., 2024) highlight the difficulty of robust multi-step planning, especially in open-ended tasks. To address this, methods like task decomposition (Zhang et al., 2024) and ReAct, which interleaves reasoning and acting, have been shown to improve agent adaptability (Yao et al., 2023), especially when used in collaborative or multi-agent domains (Lal et al., 2023).
However, fully autonomous agents like AutoGPT often struggle with robustness and transparency in real-world use (Zheng et al., 2024; Yao et al., 2024). Cognitive research suggests humans rarely plan exhaustively upfront (Krieger et al., 2009; Teevan et al., 2016), pointing to a need for agents that support incremental and context-sensitive planning. Recent studies show that exposing agent reasoning and allowing user intervention improves both performance and trust (He et al., 2025). In sum, planning in agentic AI is not just a technical hurdle; it is a design problem. Transparent reasoning, iterative refinement, and human-in-the-loop oversight are essential for building trustworthy agents that can meaningfully support real-world tasks.
@@ 2.3. Human Simulations with LLMs
Recent work has shown that LLMs in multi-agent simulations can exhibit surprisingly human-like behaviors, forming routines, organizing events, and coordinating spontaneously without explicit instructions (Park et al., 2023; Horton, 2023; Park et al., 2022). These behaviors often reflect human reasoning and social strategies, aligning with patterns observed in economic games (Sreedhar et al., 2025; Sreedhar and Chilton, 2025; Aher et al., 2023), judicial decisions (Hamilton, 2023), and over 80% of psychology experiments (Hewitt et al., 2024). This suggests LLM agents can serve as useful proxies for modeling human behavior at scale.
Beyond simulation, LLMs are demonstrating real-time social competence. CICERO, a dialogue-enabled Diplomacy agent, ranked in the top 10% of human players by integrating language with strategic planning (Team et al., 2022). Similarly, the MATRIX framework used simulated AI dialogues to surface moral tradeoffs in Constitutional AI evaluations (Pang et al., 2024), showing how LLMs can assist in value alignment. LLM simulations also support red teaming and adversarial testing, enabling safe, scalable evaluation of edge cases and vulnerabilities (Purpura et al., 2025). While imperfect, these agents provide a reproducible and low-risk method for probing system behavior in socially sensitive contexts, offering a powerful complement to human-centered alignment and evaluation.
@@ 3. DoubleAgents System
Figure 3. A screenshot of DoubleAgents coordinating the assignment of four speakers to four seminar slots. (A) Policy Panel – Displays coordination policies that reflect user values and preferences. Users can add, edit, or delete policies, which guide the reasoning and planning of the coordination agent. (B) Interactive Chat Interface – Shows the full history of plans and actions, allowing users to monitor planning and execution, confirm or modify agent suggestions, and intervene when necessary. (C) Plan and Action Pop-up – A focused view within the chat history where users can review, edit, regenerate, or approve individual plans and actions. (D) Assignment Tracker – Visualizes tentative speaker-to-slot assignments and indicates each speaker’s availability. Dark purple denotes confirmed availability; light purple signals ambiguous availability requiring user confirmation. (E) Communication History Viewer – Displays the messaging history between the coordination agent and each speaker, organized by date and interaction. (F) Calendar View – Provides temporal context by showing the current simulation and upcoming seminar dates, color-coded for clarity.
We present DoubleAgents, a system for aligning agentic AI through interactive simulation. Built on the ReAct framework (Yao et al., 2023), DoubleAgents supports iterative planning and execution via a team of AI agents that collaborate with users on complex coordination tasks. In this paper, we focus on a concrete use case: assigning seminar speakers to available time slots (see Appendix A for additional background). DoubleAgents takes as input a user-defined goal (e.g., “Organize a seminar with four speakers and four slots”), persona descriptions for simulated respondent agents, the user’s existing email templates, and a set of generic scheduling policies. Through iterative simulations, the system produces a finalized assignment of speakers to slots, along with refined policies, updated email templates, and executable stop hooks that users can later deploy in real-world agentic workflows. The system comprises six core components organized into three functional categories. First, DoubleAgents includes the Core ReAct Agents, which implement the reasoning, planning, and simulation logic that drives the coordination process:
  • Coordination Agent - Responsible for reasoning over the current problem state using a set of predefined policies. It proposes plans and actions, and invokes appropriate tools to execute them.
  • Simulated Respondents - Represent background agents that emulate human respondents’ behavior by responding to user-generated emails based on predefined personas and descriptions.
Second, DoubleAgents incorporates the System Orchestration Modules, which manage the internal context and ensure robustness during decision-making:
  • Context Management Module - Maintains all relevant contextual information related to the task, including simulated respondents’ details (e.g., personas and descriptions), email logs from prior time steps, and the history of user-approved plans, actions, and emails.
  • Edge Case Detector - Identifies situations that fall outside the coverage of existing policies and escalates them to the user for clarification and additional instructions. These escalations are recorded as stop hooks, enabling iterative refinement and extension of the policy set.
Lastly, the User Interaction Interfaces components enable transparency, user control, and real-time feedback during the coordination process:
  • Interactive Chat Interface - Facilitates communication between the user and the system. Users can approve or revise generated plans, actions, and emails, and respond to flagged edge cases by providing natural language feedback.
  • Visualization Dashboard - Provides real-time feedback on the state of the task, including temporal progress, assignment status, and an activity tracker to help users monitor the overall coordination process.

### v3

# DoubleAgents: Human-Agent Alignment in a Socially Embedded Workflow

@@ Abstract.
Aligning agentic AI with user intent is critical for delegating complex, socially embedded tasks, yet user preferences are often implicit, evolving, and difficult to specify upfront. We present DoubleAgents, a system for human–agent alignment in coordination tasks, grounded in distributed cognition. DoubleAgents integrates three components: (1) a coordination agent that maintains state and proposes plans and actions, (2) a dashboard visualization that makes the agent’s reasoning legible for user evaluation, and (3) a policy module that transforms user edits into reusable alignment artifacts, including coordination policies, email templates, and stop hooks, which improve system behavior over time. We evaluate DoubleAgents through a two-day lab study (n=10), three real-world deployments, and a technical evaluation. Participants’ comfort in offloading tasks and reliance on DoubleAgents both increased over time, correlating with the three distributed cognition components. Participants still required control at points of uncertainty — edge-case flagging and context-dependent actions. We contribute a distributed cognition approach to human-agent alignment in socially embedded tasks.
generative AI, AI agents, trust, reliance, transparency, state tracking, seminar organizing, control, human coordination, constitutional AI
copyright: acmcopyrightjournalyear: 2025conference: 2026; 2026; UISTbooktitle: Proceedings of the UIST 2026price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
Figure 1. DoubleAgents aligns agentic AI with user preferences for socially embedded coordination tasks such as organizing seminars. From an initial setup (left), including a goal, simulated personas, and seed policies and templates, users iteratively align the system (middle) by reviewing proposed plans, approving actions, handling edge cases, and refining policies, templates, and stop hooks across repeated runs. This produces (right) both a finalized speaker-slot allocation and a set of user-aligned artifacts — policies, email templates, and stop hooks — that persist and transfer to future coordination tasks.
@@ 1. Introduction
Alignment between humans and AI is the foundation of useful agentic systems. To act on a user’s behalf, an agent must understand intent that is often implicit, evolving, and context-dependent — internalizing tacit norms of when (not) to act, when (not) to ask, and how to match a user’s style without being told each time. Without alignment, small misunderstandings scale into systemic problems across workflows and social relationships, eroding the trust that makes delegation possible.
Even well-aligned AI still needs human oversight, both to correct errors and to guide new edge cases in complex, real-world settings. The end goal of alignment is not to fully replace human judgment but to let humans offload as much as possible to AI while maintaining control. How much is offloaded should be up to the human: the more aligned the system, the more confidently a user can delegate. But alignment cannot be specified upfront. Users often do not know what they want until they see the system act, and their preferences evolve as the task unfolds. The central question is how humans and AI can work together through repeated interaction and refinement to reach alignment and then comfortable offloading.
Distributed cognition (Hollan et al., 2000; Hutchins and Klausen, 1996) offers a principled framework for designing this kind of human-AI partnership. Research on complex sociotechnical systems has shown that for many hard problems, like flying an airplane, no single individual is fully responsible for the cognition. Instead, cognition is distributed across humans and artifacts: the pilot, co-pilot, autopilot, cockpit instruments, ground control, and maintenance crews all contribute. Three components make this work. First, tracking state: the system must maintain a shared sense of where things stand, as cockpit instruments monitor altitude, speed, and position in real time. Second, representing information: critical state must be displayed in the right form so actors can make judgments at a glance, as cockpit displays show system overviews rather than raw data. Third, adapting policies: all actors follow procedures and checklists that encode accumulated knowledge about safe operation, as aviation checklists evolve after incidents to cover new situations.
We apply this framework to design DoubleAgents, an interactive system for human-agent alignment in socially embedded coordination tasks. The name reflects the system’s two types of AI agents: a coordination agent that manages the task on behalf of the user, and simulated respondent agents — either real recipients in live deployment or simulated AI agents that stand in for real people during rehearsal runs. This dual mode lets users iterate on alignment through either live deployment with realistic feedback or simulation that compresses weeks into minutes before real social stakes are involved. Our domain is organizing a university speaker seminar series, a task where a human must monitor evolving situations over weeks, follow up with the nonresponsive, and handle unexpected scenarios, all while navigating real social stakes (you cannot unsend an email, nor be too pushy or too passive).
DoubleAgents distributes cognition through three mechanisms: a coordination agent that tracks all state and proposes next actions, serving as the organizer’s memory and reasoning partner; a dashboard with temporal, social, and procedural views that make the agent’s reasoning legible so the user can verify and evaluate proposals; and a policy module that turns users’ direct edits into reusable alignment artifacts (coordination policies, email templates, and stop hooks) that improve over repeated runs through either live deployment or simulation.
Overall, we make three contributions:
  • A distributed cognition framework for human–AI alignment in socially embedded tasks, identifying three core components: tracking coordination state, externalizing information through visualization, and adapting behavior via policies, and articulating how these jointly support effective alignment.
  • DoubleAgents, a system that operationalizes this framework through three components: (1) a coordination agent that maintains shared state and proposes actions, (2) a visualization layer that makes reasoning legible for user evaluation, and (3) a policy module that enables users to iteratively refine reusable alignment artifacts. The system supports both simulation for rapid iteration and live deployment for real-world coordination.
  • Empirical insights from controlled studies and real-world deployments demonstrating the distinct role of each component, policies for shaping planning, visualizations for supporting evaluation, and stop hooks for governing execution, and showing that iterative alignment process produces reusable alignment artifacts that persist across sessions.
@@ 2.1. Alignment in Human-AI Systems
As AI agents take on more complex tasks, ensuring they align with human values has become a central research challenge. Following the Bidirectional Human–AI Alignment framework (Shen et al., 2025), we view alignment in agentic settings as an ongoing co-adaptation process. On the system side, agents must be continuously shaped through interaction history and curated context so their actions reflect a user’s goals, preferences, and social norms (Zhang et al., 2025; Gao et al., 2024). On the human side, users must develop appropriate mental models and usage patterns to effectively direct, evaluate, and calibrate delegation (Vanderlyn et al., 2025; Andrews et al., 2023). Operationalizing this bidirectional view requires mechanisms that let users specify intent, oversee execution, and revise expectations over time. Terry et al. (2023) decomposes this challenge into specification, process, and evaluation alignment, capturing the full arc from goal communication through output verification. Subsequent work has addressed each stage: Shankar et al. (2024) shows with EvalGen that users’ evaluation criteria themselves drift as they observe model outputs, while Epperson et al. (2025) supports process-level steering with AGDebugger, enabling what-if probing and checkpoint-based resets to diagnose misalignment in multi-agent workflows. Together, these efforts outline a design space where guardrails, prompt optimization, and interactive debugging serve as complementary post-hoc alignment mechanisms.
@@
Constitutional AI offers a different strategy, replacing constant human feedback with rule-based self-supervision using explicit principles (Bai et al., 2022). This has been expanded into participatory frameworks such as Collective Constitutional AI, where large groups co-author alignment rules to reflect diverse values (Huang et al., 2024), and ConstitutionMaker, which helps users turn natural language feedback into usable policies (Petridis et al., 2024). DoubleAgents contributes to this space by embedding alignment into an interactive process. Following Suchman’s view of plans as situated resources rather than fixed programs (Suchman, 1987), policies and stop hooks in our system function as evolving artifacts that users negotiate in context. Through interactions, users iteratively refine coordination policies, email tone, and stop conditions, treating alignment as a continuous, participatory process shaped by the situations in which it unfolds.
2.2. Agentic AI: From Task Planning to Socially Embedded Workflows
Delegating complex tasks to AI agents raises concerns around trust, transparency, and alignment (Afroogh et al., 2024), motivating designs that keep humans in the loop. Exposing agent reasoning with user intervention improves both performance and trust (He et al., 2025), yet cognitive research suggests humans rarely plan exhaustively upfront (Krieger et al., 2009; Teevan et al., 2016), pointing to a need for agents that support incremental, context-sensitive planning with human oversight. Whether agents operate through tool use (Schick et al., 2023) or direct interface manipulation (Anthropic, 2025; OpenAI, 2025; Manus AI, 2025; Labs, 2025), the difficulty of robust multi-step planning has been consistently highlighted in recent benchmark works (Yao et al., 2024; Zheng et al., 2024; Xie et al., 2024; Trivedi et al., 2024), underscoring that oversight is not merely desirable but necessary.
These challenges intensify when agents operate in socially embedded contexts, where correctness depends on interpersonal norms and situated judgment rather than functional specifications. Benchmarks for social intelligence reveal a wide capability gap: SOTOPIA (Zhou et al., 2024) found that LLM agents significantly underperform humans on social commonsense, while CASA (Qiu et al., 2025) showed high norm violation rates in agentic settings, suggesting task-directed focus actively crowds out social sensitivity. DoubleAgents targets the gap this literature exposes: the absence of end-to-end systems for multi-turn, socially embedded coordination. Rather than maximizing autonomy, our system helps users iteratively author and refine agent behavior through calibrating delegation boundaries, coordination policies, and communication tone, for tasks where social context and timing are integral to correctness.
@@ 2.3. Human Simulations with LLMs
Recent work has shown that LLM-based multi-agent simulations can exhibit human-like behaviors in social and coordination settings, forming routines, organizing events, and interacting in ways that resemble human reasoning patterns (Park et al., 2023; Horton, 2023; Park et al., 2022). These simulations have been used to approximate human behavior in domains such as economic decision-making (Sreedhar et al., 2025; Sreedhar and Chilton, 2025; Aher et al., 2023), judicial reasoning (Hamilton, 2023), and psychology experiments (Hewitt et al., 2024). They have also supported applications such as strategic gameplay (Team et al., 2022), adversarial testing (Purpura et al., 2025), and value alignment through simulated interactions (Pang et al., 2024).
In DoubleAgents, LLM-simulated respondents are used to provide a fast and flexible way to surface a range of coordination scenarios, including potential edge cases, so that users can iteratively refine their delegation strategies. These interactions primarily serve to accelerate the alignment process by making system behavior observable early, while similar workflows can also be carried out during real-world deployment with actual participants.
@@ 3. DoubleAgents System
Figure 2. A snapshot of DoubleAgents coordinating four speakers across seminar slots. Grounded in distributed cognition, the system organizes interaction across three components: (a) Policies - encoding evolving user preferences and procedures; (b) The Coordination Agent - tracking state, plans next steps, suggests actions, and surfaces the policies guiding its decisions; and (c) The Dashboard Visualization - externalizing key information, including assignments, timelines, and communication history, enabling users to quickly interpret the situation. In use, the coordination agent monitors progress, proposes a plan, and explains its rationale; the user reviews the state through the dashboard, verifies that the plan aligns with expectations, and approves it to proceed. Through interactions, policies are refined, allowing the system to better align with the user over time.
DoubleAgents is an interactive system for aligning agentic AI behavior with user preferences in coordination tasks. We ground the system in a concrete use case: organizing a university speaker seminar series (see Appendix A for additional background), a setting where users must manage evolving constraints, follow up with nonresponsive participants, and handle unexpected situations under real social stakes. Consistent with our distributed cognition framework, DoubleAgents distributes coordination across three components: a coordination agent that tracks task state and proposes next plans and actions, a dashboard that presents temporal, social, and procedural views to make the agent’s reasoning legible for user evaluation, and a policy module that turns user edits into reusable alignment artifacts (e.g., coordination policies, email templates, and stop hooks). Together, these components support iterative refinement of agent behavior over time, enabling users to progressively align the system through interaction in both simulated and real-world deployment settings. In this section, we first present a user walkthrough (Section 3.1), followed by the system implementation (Section 3.2).
3.1. User Walkthrough